Circular No. 8013

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2002 V2 (LINEAR) An apparently asteroidal object discovered by LINEAR (discovery ob-

servation below) and posted on the NEO Confirmation Page has been found to have cometary appearance. M. Tichý found a compact 6′′ coma on KLENOT images taken at Kleť on Nov. 7.13 UT. Images taken by P. R. Holvorcem and M. Schwartz with the Tenagra 0.81-m reflector on Nov. 7.4 and 8.4 show the object diffuse with a coma diameter of ∼ 8′′.

2002 UT  α_{2000}  δ_{2000}  m_{2}
Nov. 5.31161  5^{h}41^{m}06.^{s}16  +10^{o}31′54′′.8  18.6

Additional astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2002-V48.

T = 2003 Jan. 10.559 TT  ω = 304.^{o}195
Ω = 19.392 }2000.0
q = 6.94322 AU  i = 166.881

SUPERNOVAE 2002gx, 2002hu, AND 2002hv T. Matheson, P. Challis, and R. Kirshner, Harvard-Smithsonian Center

for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2002hu (cf. IAUC 8012), obtained by L. Macri on Nov. 8.80 UT with the Mt. Hopkins 1.5-m telescope (+ FAST), shows it to be a type-Ia supernova before maximum. Narrow Hα emission from a nearby H II region indicates a recession velocity of ∼ 11000 km/s for the host galaxy. The expansion velocity derived from the minimum of the Si II line (rest 635.5 nm) is ∼ 10500 km/s. The spectral-feature age of the supernova is 5 ± 2 days before maximum light. A spectrum of SN 2002hv (cf. IAUC 8012), obtained by P. Berlind on Nov. 9.49 with the same telescope, shows it to be a type-Ia supernova a few weeks past maximum.

A. V. Filippenko, R. Chornock, and S. Jha, University of California, Berkeley; and B. Barris, University of Hawaii, report that inspection of CCD spectra (range 310–1000 nm), obtained on Nov. 8 UT with the Keck I 10-m telescope (+ LRIS), finds that SN 2002gx (IAUC 7995) is of type Ia, ∼ 2 months past maximum brightness. The redshift of the host galaxy is roughly 0.08. SN 2002hu is also of type Ia, a few days before maximum brightness. Once again, from a comparison of supernova spectra, the redshift of the host galaxy is ∼ 0.03. SN 2002hv is also of type Ia, ∼ 2 months past maximum.

2002 November 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green